a= root 3 +(inside) root 3 + inside root 3...
what is the method to solve such questions
 

Grade:10

1 Answers

sonika p
6 years ago
Since it`s an infinite series, a = root 3 +(inside) root 3 + inside root 3... which is similar to a = root (3 + a) squaring on both sides, a^2 = 3 + aa^2 -a -3 = 0 solving the quadratic equation,